OpenStudy (anonymous):

Help will give medal Factor: 12s^2-23s+5

OpenStudy (anonymous):

Use the Factoring AC method. Find 2 numbers b1 and b2 that: b1*b2 = 60, and (b1 + b2) = -23. b1 = -3 and b2 = -20. Replace in the trinomial the term (-23x) by the 2 terms (-3x) and (-20x), then factor.

OpenStudy (ybarrap):

We pretty much did this one just now... $$ 12 s^2-23 s+5 = 12 s^2-20 s-3 s+5 = 3 s (4 s-1)-5 (4 s-1) $$ http://openstudy.com/study#/updates/5326485ce4b0ba8d4c415290

OpenStudy (anonymous):

= (4s -1)(3s - 5)